OpenStudy (anonymous):

In an infinite geometric sequence, what is the sum in this pattern? 8+4+2+1+...

OpenStudy (anonymous):

Well, 1+.5+.25+... = 2 So 8+4+2+2

OpenStudy (anonymous):

sorry i dont get it but thanks for trying to help!

OpenStudy (anonymous):

\(\frac{4}{8}=\frac{1}{2}\) so you have a geometric series with \(r=\frac{1}{2}\) the general formula is \[\frac{a}{1-r}\] where \(a\) is the first term in your example \(a=8\)

OpenStudy (anonymous):

you can do this in your head what is one minus one half?

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

oh omg

OpenStudy (anonymous):

thanks LOL i wasn't thinking......

OpenStudy (anonymous):

wait you are not done one minus one half is one half what is the reciprocal of one half?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

i already got the answer by dividing 8/0.5 though